OpenStudy (anonymous):

Which statement is correct about a sphere? It has exactly one great circle. Its great circle has the same radius as the sphere. The areas of its great circles through different planes are not equal. The diameter of its great circle is double the diameter of the sphere.

ganeshie8 (ganeshie8):

any plane passing thru center creates a great circle. so a sphere has infinite great circles.

ganeshie8 (ganeshie8):

so you can strike off first option

OpenStudy (anonymous):

Okay, I'm going to guess B?

ganeshie8 (ganeshie8):

and thats a very good guess :)

ganeshie8 (ganeshie8):

great circle radius = sphere radius

ganeshie8 (ganeshie8):

thats the oly correct option. last two options are incorrect.

OpenStudy (anonymous):

Awesome, could you help me with some more?

ganeshie8 (ganeshie8):

sure :)

OpenStudy (anonymous):

i would choose second

OpenStudy (anonymous):

A company that manufactures storage bins for grains made a drawing of a silo. The silo has a conical base, as shown below. What is the total volume of grains that can be stored in the silo? 34π ft3 68π ft3 136π ft3 152π ft3

ganeshie8 (ganeshie8):

lol i did this q like 20 times this year

OpenStudy (anonymous):

Lol I've been working on it for like 2 days

ganeshie8 (ganeshie8):

First, see that its a mix of two shapes : cone, cylinder

ganeshie8 (ganeshie8):

both have same radius, r = 4/2 = 2

ganeshie8 (ganeshie8):

cylinder has height h = 8

ganeshie8 (ganeshie8):

cone has height h = 9.5 - 8 = 1.5

OpenStudy (anonymous):

Right, I got that yesterday but somehow didn't get the right answer at the end

ganeshie8 (ganeshie8):

volime of silo = volume of cylinder + volume of cone = \(\pi r^2 h + \frac{1}{3}\pi r^2 h\) = \(\pi \ 2^2 8 + \frac{1}{3}\pi \ 2^2 1.5\) = \(32 \pi + 2\pi \) = \(34 \pi \)

ganeshie8 (ganeshie8):

see if it makes some sense :)

OpenStudy (anonymous):

Yeah actually it does, lol thank you! You have time for more? (:

ganeshie8 (ganeshie8):

yea sure

OpenStudy (anonymous):

Just tell me when you're sick of me haha

ganeshie8 (ganeshie8):

lol sure you're awesome carry on :)

OpenStudy (anonymous):

The figure below shows the ideal pattern of movement of a herd of cattle, with the arrows showing the movement of the handler as he moves the herd. The arc the handler makes from the starting point to the return point should be a quarter of a circle. Based on this theory, what distance will the handler move from the starting point to the return point if he creates an arc of a circle of radius 70 feet?

ganeshie8 (ganeshie8):

easy, use perimeter of circle formula

ganeshie8 (ganeshie8):

r = 70

ganeshie8 (ganeshie8):

perimeter = \(2 \pi r\) = ?

ganeshie8 (ganeshie8):

wat ever u get, u need to divide it by 4, since it is a quarter circle.

OpenStudy (anonymous):

What do I do first? lol

ganeshie8 (ganeshie8):

plugin r = 70 in above formula

OpenStudy (anonymous):

Right, but then what do i do first in that formula

ganeshie8 (ganeshie8):

perimeter = \(2\pi r\) = \(2 \pi . 70\) = \(140 \pi\)

ganeshie8 (ganeshie8):

we need to divide it by 4 to get the quarter circle perimeter

ganeshie8 (ganeshie8):

perimeter of quarter circle = \(\frac{140 \pi}{4}\) = ?

OpenStudy (anonymous):

So 109.9?

ganeshie8 (ganeshie8):

you may want to leave it exact, instead of converting it to decimals

ganeshie8 (ganeshie8):

\(\frac{140 \pi}{4}\) \(35 \pi\)

OpenStudy (anonymous):

Nah my answers are in decimal form lol

OpenStudy (anonymous):

@ganeshie8 A cylindrical piece of iron pipe is shown below. The wall of the pipe is 0.75 inch thick, and the pipe is open at both ends. What is the approximate inside volume of the pipe?

ganeshie8 (ganeshie8):

lol ok then you're right :)

OpenStudy (anonymous):

Alrighty, what about the one I just posted?

ganeshie8 (ganeshie8):

ok, id let u do this.. give it a try :)

ganeshie8 (ganeshie8):

Given, total diameter = 6

ganeshie8 (ganeshie8):

within that, how much goes waste for thickness ?

OpenStudy (anonymous):

5.25?

ganeshie8 (ganeshie8):

careful, thickness is on both sides of diameter !

ganeshie8 (ganeshie8):

so you need to remove 0.75 two times

OpenStudy (anonymous):

Okay so 4.50?

ganeshie8 (ganeshie8):

thats diameter of inner space.

ganeshie8 (ganeshie8):

wats the radius ?

OpenStudy (anonymous):

2.25

ganeshie8 (ganeshie8):

Correct ! you have everything u need to use the formula : volume of cylinder = \(\pi r^2 h\) \(r = 2.25\) \(h = 16\)

ganeshie8 (ganeshie8):

plugin

OpenStudy (anonymous):

Alright so 254.34?

ganeshie8 (ganeshie8):

you have used \(\pi = 3.14\) is it

ganeshie8 (ganeshie8):

then yes you're right :D

OpenStudy (anonymous):

It's easy once I know what to do lol. How would I find the volume of this one? A cylindrical vase has a diameter of 6 inches. At the bottom of the vase, there are 9 marbles, each of diameter 2 inches. The vase is filled with water up to a height of 12 inches. What is the volume of water in the vase?

ganeshie8 (ganeshie8):

two steps

ganeshie8 (ganeshie8):

step 1 : find the volume of marbles

ganeshie8 (ganeshie8):

step 2 : find the volume of water including marbles

ganeshie8 (ganeshie8):

step 3 : subtract 1 from 2

ganeshie8 (ganeshie8):

ok those are 3 steps lol

ganeshie8 (ganeshie8):

... 9 marbles, each of diameter 2 inches.

ganeshie8 (ganeshie8):

diameter = 2, so r = 1

ganeshie8 (ganeshie8):

volume of one marble = \(\frac{4}{3} \pi r^3 \) = \(\frac{4}{3} \pi 1^3 \) = \(\frac{4}{3} \pi \)

ganeshie8 (ganeshie8):

volume of 9 marbles = \(9 \times \frac{4}{3} \pi \)\) = \(12 \pi\) -----------------------(1)

ganeshie8 (ganeshie8):

you wid me so far ha

OpenStudy (anonymous):

I'm trying lol

ganeshie8 (ganeshie8):

ok we're done wid step1. lets move to step2

ganeshie8 (ganeshie8):

The vase is filled with water up to a height of 12 inches. So, volume of water including marbles = \(\pi r^2 h\) = \(\pi 3^2 12\) = \(108 \pi \) ------------------(2)

ganeshie8 (ganeshie8):

volume of water alone = (2) - (1) = \(108\pi - 12\pi\) = ?

OpenStudy (anonymous):

376.80

ganeshie8 (ganeshie8):

check again

OpenStudy (anonymous):

96?